Levie tosses an unbiased coin. He receives PHP 100 if a head appears and he pays PHP 40 is a tail appears. Find the expected value and the variance of his gain.


1
Expert's answer
2021-06-23T06:15:40-0400

Let "X=" the amount of money which Levie receives.


"\begin{matrix}
 x & 100 & -40 \\
 p(x) & 0.5 & 0.5
\end{matrix}"

"E[X]=0.5(100)+0.5(-40)=30"

"E[X^2]=0.5(100)^2+0.5(-40)^2=5800"

"V(X)=E[X^2]-(E[X])^2"

"=5800-(30)^2=4900"